(-2,16)(0,4)(4,28) make into standard form using parabola
Answer by ankor@dixie-net.com(22740)

(-2,16)(0,4)(4,28) make into standard form using parabola
:
Using the form ax^2 + bx + c = y
:
The 2nd coordinate; 0,4, tells us the y intercept therefore: c = 4
:
Use the 1st and 3rd coordinates to find a and b
x=-2; y=16
-2^2a - 2b + 4 = 16
4a - 2b = 16 - 4
4a - 2b = 12
simplify, divide 2
2a - b = 6
:
x=4, 7=28
4^2a + 4b + 4 = 28
16a + 4b = 28 - 4
16a + 4b = 24
Simplify, divide by 4
4a + b = 6
:
Add the two simplified equations
2a - b = 6
4a + b = 6
---------------adding eliminates b, find a
6a = 12
a = 2
:
Find b using the 1st simplified equation
2a - b = 6
2(2) - b = 6
-b = 6 - 4
-b = 2
b = -2
;
The equation: y = 2x^2 - 2x + 4; in standard form